# Multi-Level Explanations for Generative Language Models

### v1

@@ Appendix B Algorithm Parameter Settings
We set the number of model queries (parameter max_evals) to be approximately equal to the number used by our L-SHAP and C-LIME algorithms. More specifically, since obtaining mixed-level attributions with MExGen requires first performing sentence-level attribution, we add the numbers of queries used during sentence-level and mixed-level attribution. We then take the larger of these two sums for L-SHAP and C-LIME as the number of queries allowed for PartitionSHAP.
Appendix C Automated Evaluation Details and More Quantitative Results
Computing environment
Experiments were run on a computing cluster providing V100 and A100 GPUs with 32 GB of CPU memory and 32, 40, or occasionally 80 GB of GPU memory. One CPU and one GPU were used at a time. The total computation time is estimated to be on the order of 100–200 hours.
Datasets
For summarization, we limit our evaluation to the first 100 examples from the test sets of XSUM and CNN/Daily Mail, due to the computation involved in running multiple perturbation-based attribution algorithms that query summarization models. For question answering on SQuAD, we evaluate 200 randomly sampled examples from the validation set.
Perturbation curves
MExGen can attribute to mixed units of different lengths in terms of the number of tokens, and these units also differ from those produced by P-SHAP. To account for these differences in computing perturbation curves, we consider both the attribution score and the number of tokens for each unit. We rank units in decreasing order of the unit attribution score divided by the number of tokens (since we plot perturbation curves as functions of the number of tokens perturbed, this ratio can be seen as the slope in the score-tokens plane). We then perturb (more precisely remove) the top $k$ units according to these ratios, compute the output score given by the evaluation scalarizer, and increase $k$ until at least $50$% of the tokens have been removed. To average perturbation curves over examples with different numbers of tokens, we divide the numbers of tokens perturbed by the total number to obtain percentages. We then linearly interpolate onto a common grid of percentages before averaging.
@@ Appendix C Automated Evaluation Details and More Quantitative Results
Figure 6 shows the cosine similarity between all the pairs of secalarizers we use. We show the cosine similarity between the explanations of each method to analyze how aligned the explanations are from different scalarizers. Each text input $x=x_{1},...,x_{d}$ receives a multi-level explanation given by $\xi_{S}(x)=(\xi_{S}(x)_{1},...,\xi_{S}(x)_{d})\in\mathbb{R}^{d}$ where each $\xi_{S}(x)_{i}$ represents the contribution of unit $i$ to the model prediction scalarization computed using the scalarizer $S$. We define the cosine similarity between scalarizers $S$ and $S^{\prime}$ as the average of the cosine similarities between the explanations for all available input texts, i.e.,
@@
[…]
Similarity to Logist. In the SHAP library, SUMM is proposed to provide explanations to LLMs that do not provide access to the logits — hence, the main objective of SUMM is to approximate the explanations for the logit when it is not available. However, Figure 6 shows that SUMM is not always the best scalarizer for approximating the explanations that would be given if logits were available. For example, Figure 6 (a) shows that the similarity between SUMM and logit is near $0.61$. In contrast, Figure 6 (b) shows that the similarity between the explanations generated using SUMM as scalarizer has a similarity of $0.78$ with the onex generated using logits.
@@
Figures 7, 8, 9 show the perturbation curve for different scalarizations.
(a) BERT Score as Evaluation
(b) Log Prob as Evaluation
(c) BART Score as Evaluation
Figure 7: Perturbation curves for MExGen L-SHAP with different scalarizers, used to explain the DistilBart model on the XSUM dataset. The curves show the decrease in (a) BERTScore, (b) log probability, and (c) BARTScore when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error.
(a) BERT Score as Evaluation
(b) Log Prob as Evaluation
(c) BART Score as Evaluation
Figure 8: Perturbation curves for MExGen C-LIME with different scalarizers, used to explain the Flan-T5-Large model on the SQUaD dataset. The curves show the decrease in (a) BERTScore, (b) log probability, and (c) BARTScore when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error.
(a) BERT Score as Evaluation
(b) Log Prob as Evaluation
(c) BART Score as Evaluation
Figure 9: Perturbation curves for MExGen L-SHAP with different scalarizers, used to explain the Flan-T5-Large model on the SQUaD dataset. The curves show the decrease in (a) BERTScore, (b) log probability, and (c) BARTScore when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error.
Appendix D User Study
D.1 Participants
@@
D.2 Scalarizers
@@ D.3 Attribution Methods
D.4 Granularity Preference
@@
D.5 Survey
Figure 10 and Figure 11 show some questions we asked in the survey.
@@ D.5 Survey
Appendix E Future Directions
@@
It could be profitable in future work to incorporate the hierarchical explanations discussed in Section 2 into the multi-level MExGen framework. The method of Chen and Jordan [3] may be especially relevant since it leverages a constituency parse tree to compute word-level importances, which may be related to our use of dependency parse trees.
@@
We have explored perturbing words by masking them and then calling a BERT model to fill the masks with different words that fit within the sentence. However, we have thus far not seen a quantifiable benefit to using BERT compared to replacing with a fixed baseline value (such as an empty string). Our experience is in line with the the mixed results reported by Pham et al. [20] on using BERT in this manner.

### v2

@@ C.4 Baseline attribution methods
We set the number of model queries (parameter max_evals) to be approximately equal to the number used by our L-SHAP and C-LIME algorithms. More specifically, since obtaining mixed-level attributions with MExGen requires first performing sentence-level attribution, we add the numbers of queries used during sentence-level and mixed-level attribution. We then take the larger of these two sums for L-SHAP and C-LIME as the number of queries allowed for PartitionSHAP. All other parameters were kept at their default values in the SHAP library. [^12] https://github.com/shap/shap, made available under the MIT license
CaptumLIME
The number of model queries was also set as described above for PartitionSHAP. As discussed in Section 4.1, the units for attribution are the same as used by MExGen C-LIME. All other parameters were kept at their default values in the Captum library. [^13] https://github.com/pytorch/captum, distributed under the BSD 3-Clause license
@@
LLM Self-Explanation
Figure 6 shows the prompt used to produce LLM self-explanations for the experiment in Section 4.4, where the task is summarization. We first repeat the LLM’s summary back to it. We then segment the input article into the same units units[0], units[1], …, (sentence-level or mixed-level) used by MExGen. This controls for input segmentation as done in the comparison with CaptumLIME. To make the ranking task easier, we followed LongBench-Cite (Zhang et al., 2024) in prepending a numbered tag (<u0>, <u1>, …) to each unit. The LLM is asked to list only the tags corresponding to units, in order of decreasing importance for generating the summary. The number top_k of units to be ranked is set to $30\%$ of the total number of units. Since we evaluate perturbation curves up to $20\%$ of the total tokens, the $30\%$ setting provides some margin while not requiring the LLM to rank all units.
The comma-separated output from the LLM is parsed into a list of unit indices. Elements that do not yield an integer or yield an out-of-range integer are dropped (this happened rarely however).
C.5 Perturbation curves
MExGen can attribute to mixed units of different lengths in terms of the number of tokens, and these units also differ from those produced by P-SHAP. To account for these differences in computing perturbation curves, we consider both the attribution score and the number of tokens for each unit. We rank units in decreasing order of the attribution score divided by the number of tokens (since we plot perturbation curves as functions of tokens perturbed, this ratio can be seen as the slope in the score-tokens plane). Thus, a shorter unit is ranked higher than a longer unit if both have the same attribution score. We then perturb (more precisely remove) the top $k$ units, compute the output score given by the evaluation scalarizer, and increase $k$ until at least $20$% of the tokens have been removed.
To average perturbation curves over examples, which have different numbers of tokens, we divide the numbers of tokens perturbed by the total number of tokens in each example to obtain percentages of tokens. We then linearly interpolate onto a common grid of percentages before averaging and computing standard errors.
C.6 Computing environment
Experiments were run on a computing cluster providing nodes with 32 GB of CPU memory, V100 GPUs with 32 GB of GPU memory, and occasionally A100 GPUs with 40 or 80 GB of GPU memory. One CPU and one GPU were used at a time. The total computation time is estimated to be on the order of 1000 hours.
Appendix D More Automated Evaluation Results
D.1 Scalarizer Evaluation
@@
Figure 7 shows the cosine similarity between all the pairs of scalarizers we use. We show the cosine similarity between the explanations of each method to analyze how aligned the explanations are from different scalarizers. Each text input $x=x_{1},...,x_{d}$ receives a multi-level explanation given by $\xi_{S}(x)=(\xi_{S}(x)_{1},...,\xi_{S}(x)_{d})\in\mathbb{R}^{d}$ where each $\xi_{S}(x)_{i}$ represents the contribution of unit $i$ to the model prediction scalarization computed using the scalarizer $S$. We define the cosine similarity between scalarizers $S$ and $S^{\prime}$ as the average of the cosine similarities between the explanations for all available input texts, i.e.,
@@
(a) Log Prob as Evaluation
(b) BERT Score as Evaluation
(c) SUMM Score as Evaluation
Figure 8: Perturbation curves for MExGen L-SHAP with different scalarizers, used to explain the DistilBART model on the XSUM dataset. The curves show the decrease in (a) log probability, (b) BERTScore, and (c) SUMM score when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error.
@@
Similarity Across Scalarizers. Figure 7 indicates that, although some scalarizers lead to similar model explanations, there are occasions where scalarizers are more dissimilar. Moreover, the similarities between scalarizers not only depend on the scalarizer itself but also on the model and dataset being explained. For example, Figure 7 (b) shows that when using MExGen L-SHAP to provide explanations for the predictions in the XSUM dataset using DistilBrat, the scalarizers BERT and SUMM are fairly similar ($\texttt{CosSim}(S_{\textit{BERT}},S_{\textit{SUMM}})=0.96$). On the other hand, looking at the same pair of scalarizers but for MExGen LIME to provide explanations to the predictions in the SQUaD dataset using Flan-T5-Large, BERT and SUMM are more dissimilar ($\texttt{CosSim}(S_{\textit{BERT}},S_{\textit{SUMM}})=0.82$). This result highlights the necessity of exploring different scalarizers to explain natural language generation, taking into account the task being performed and the main objective of the explanation, i.e., target scalarization.
Similarity to Logits. In the SHAP library, SUMM is proposed to provide explanations to LLMs that do not provide access to the logits — hence, the main objective of SUMM is to approximate the explanations for the logit when it is not available. However, Figure 7 shows that SUMM is not always the best scalarizer for approximating the explanations that would be given if logits were available. For example, Figure 7 (a) shows that the similarity between SUMM and logit is near $0.61$. In contrast, Figure 7 (b) shows that the similarity between the explanations generated using SUMM as scalarizer has a similarity of $0.78$ with the onex generated using logits.
@@
(a) Log Prob as Evaluation
(b) BERT Score as Evaluation
(c) SUMM Score as Evaluation
Figure 9: Perturbation curves for MExGen C-LIME with different scalarizers, used to explain the Flan-T5-Large model on the SQuAD dataset. The curves show the decrease in (a) log probability, (b) BERTScore, and (c) SUMM score when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error.
(a) Log Prob as Evaluation
(b) BERT Score as Evaluation
(c) SUMM Score as Evaluation
Figure 10: Perturbation curves for MExGen L-SHAP with different scalarizers, used to explain the Flan-T5-Large model on the SQuAD dataset. The curves show the decrease in (a) log probability, (b) BERTScore, and (c) SUMM score when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error.
@@
Figures 8, 9, 10 show the perturbation curve for different scalarizations.
(a) Flan-UL2 on XSUM
(b) Flan-UL2 on CNN/DM
(c) Flan-T5-Large on SQuAD
(d) Llama-3 on XSUM
(e) Llama-3 on CNN/DM
(f) Llama-3 on SQuAD
Figure 11: Perturbation curves (higher is better) from different explanation methods for additional models and datasets: (a) Flan-UL2 on XSUM, (b) Flan-UL2 on CNN/DM, (d) Llama-3-8B-Instruct on XSUM, (f) Llama-3-8B-Instruct on SQuAD, plus (c) Flan-T5-Large on SQuAD and (e) Llama-3-8B-Instruct on CNN/DM repeated from Figure 5 to facilitate comparison. Shading shows standard error in the mean.
D.2 Comparison Between Explainers
Figure 11 compares the perturbation curves from MExGen instantiations and P-SHAP for additional model-dataset pairs. The patterns are similar to those in Figure 5. The one exception is in Figure 11b where P-SHAP attains a larger SUMM decrease as more tokens are perturbed, but the MExGen curves are still higher for the top 5% of tokens.
A possible reason for why P-SHAP performs better after a certain fraction of tokens in Figures 11b, 11c is as follows: P-SHAP perturbs larger subsets of the input than MExGen, for which we intentionally limit the size of perturbed subsets. These larger subsets may enable P-SHAP to find larger changes in output (higher perturbation curve) at larger fractions of tokens perturbed.
D.3 Comparison with LLM Self-Explanation
(a) Granite-3.3 on XSUM with Prob
(b) Granite-3.3 on XSUM with Prob
(c) DeepSeek-V3 on XSUM with BART
(d) Granite-3.3 on CNN/DM with Prob
(e) Granite-3.3 on CNN/DM with Prob
(f) DeepSeek-V3 on CNN/DM with BART
Figure 12: Perturbation curves comparing MExGen variants to LLM self-explanation. LLMs, datasets, and scalarizers (the latter used for both explanation and evaluation): (a) Granite-3.3-8B-Instruct on XSUM with Prob scalarizer, (b) Granite-3.3-8B-Instruct on XSUM with BART scalarizer, (c) DeepSeek-V3 on XSUM with BART scalarizer, (d) Granite-3.3-8B-Instruct on CNN/DM with Prob scalarizer, (e) Granite-3.3-8B-Instruct on CNN/DM with BART scalarizer, (f) DeepSeek-V3 on CNN/DM with BART scalarizer. The legend in (bow) applies to all panels. Shading shows standard error in the mean.
Figure 12 shows the perturbation curves corresponding to the AUPC values reported in Table 2. We also plot the perturbation curve for the intermediate sentence-level attributions (labelled “MExGen LOO (sent)”) used to obtain the mixed-level MExGen attributions, as well as the curve for the self-explanation using the same sentence-level units (“self (sent)”). The curves closely reflect the AUPC comparison already seen in Table 2. For example, MExGen greatly outperforms self-explanation when using the Log Prob scalarizer in Figure 12a, 12d. On the other hand, for the text-only BART scalarizer and larger DeepSeek-V3 model in Figure 12f, the gap is closed at higher perturbed fractions, but a small gap remains at lower fractions.
Appendix E User Study
In this section, we describe our user study.
E.1 Participants
@@
Participation in this study did not involve any significant risks beyond those present in daily life, which we explained in the consent form. At the beginning of the study, all participants read about the study procedure, risks, benefits, compensation, and costs, and provided informed consent. They voluntarily participated in the study and were free to withdraw their consent and discontinue participation at any time. Although a formal IRB process does not exist in our institution, we went through an equivalent informal process, including reviewing the study with our peers, and treated participants in accordance with ethical guidelines for human subjects.
The study was expected to take about 30 minutes or less. For compensation, each participant received 50 points (a digital currency used within the institution), which was equivalent to \$12.5 USD.
E.2 Scalarizers
@@ E.3 Attribution Methods
Figure 13: A scale used to color-code the attribution scores.
E.4 Granularity Preference
@@
E.5 Survey
The survey was structured as below with the following instructions and questions:
  1. 1.
    Consent
    • Select ’I agree’ if you are eligible and agree to the terms above. By selecting ’I agree’, you give consent to [Institution Name] to use your anonymized responses for research and development purposes. You also agree not to provide any information that is confidential or proprietary.
  2. 2.
    Input document (1): Suppose you wanted to summarize an input document and used a language model to generate a summary. Please read the text below and answer the following questions. [A randomly selected input document and a summary were inserted here]
    • Rate the overall quality of the summary. A good summary should be coherent, consistent, fluent, relevant, and accurate. (1: Poor - 7: Excellent)
  3. 3.
    Scale: Introducing the scale that was used to annotate the attribution scores. See Fig. 13
  4. 4.
    Scalarizer (2 pairwise comparisons): Using the selected ‘input document (1)’, same instructions and questions were used as shown in Fig. 14.
  5. 5.
    Input document (2): Suppose you wanted to summarize an input document and used a language model to generate a summary. Please read the text below and answer the following questions. [Another randomly selected input document and a summary were inserted here]
    • Rate the overall quality of the summary. A good summary should be coherent, consistent, fluent, relevant, and accurate. (1: Poor - 7: Excellent)
  6. 6.
    Attribution algorithms (3 pairwise comparisons): Using the selected ‘input document (2)’, same instructions and questions were used as shown in Fig. 14.
  7. 7.
    Granularity: The attribution scores can be presented in two levels of granularity, which are sentence- level and mixed-level granularities.
    • Sentence-level granularity: each sentence in the input document is color coded based on how much of it was used by the language model.
    • Mixed-level granularity: each phrase within a few high-scoring sentences is color coded based on how much of it was used by the language model. Low-scoring sentences are coded in the sentence-level granularity.
    In the next two pages, you will select your preferred granularity for each of the following examples.
    • Which granularity do you prefer in this example? [A randomly selected summary and two highlighted input documents] See Figure 15.
    • Which granularity do you prefer in this example? [Another randomly selected summary and two highlighted input documents]. See Figure 15.
  8. 8.
    Background
    • What is your job title?
    • Where is your work location?
    • What type of work do you do? (Please check all that apply.)
    • What is your proficiency level in English?
    • How often do you use language models either as part of your job or as a hobby?
    • What kind of tasks do you usually do with language models?
    • Besides the attribution scores/highlights that are shown in this study, what other information would you like to know to help you understand how a language model generated a summary?
Figure 14 and Figure 15 show primary questions we asked in the survey with screenshots.
Appendix F Future Directions
@@
It could be profitable in future work to incorporate the hierarchical explanations discussed in Appendix A into the multi-level MExGen framework. The method of Chen and Jordan (2020) may be especially relevant since it leverages a constituency parse tree to compute word-level importances, which may be related to our use of dependency parse trees.
@@
We have explored perturbing words by masking them and then calling a BERT model to fill the masks with different words that fit within the sentence. However, we have thus far not seen a quantifiable benefit to using BERT compared to replacing with a fixed baseline value (such as an empty string). Our experience is in line with the the mixed results reported by Pham et al. (2022) on using BERT in this manner.